OpenStudy (anonymous):

helpppp Solve -3x + 2 = 23 for x. (Points : 1) -7 7 -6 6 10. Solve x/5 + 4 = -2 for x. (Points : 1) 20 -20 30 -30 11. Solve 2x - 6 = 10 for x. (Points : 1) 2 4 6 8

OpenStudy (anonymous):

First you want to isolate the variable (in these problems, that would be x). To do this, add, subtract, multiply, or divide both sides (must be both sides). For the first problem you would subtract 2 from both sides (-3x = 21). Then simply divide by -3 (negative 3) and you get -7 (-7 * -3 = 21). 10. Solve x/5 + 4 = -2 Get rid of the four (so subtract four from both sides) to get x/5 = -6 then you want to get rid of the fraction (x/5) so multiply both sides by 5 to get x = -30. 11. Solve 2x - 6 = 10 Get rid of the negative 6 (so add 6 to both sides) to get 2x = 16. Then get rid of the 2 by dividing by 2 to get x = 8. Once you get the hang of it, you should be able to do it in your head. :)